An electron is released from rest in a weak electric field given by Ể--1.60 x 10-10 N/Cj. After the electron has traveled a vertical distance of 1.2 μm, what is its speed? (Do not neglect the gravitational force on the electron.) 7.65 mm/s eBook
Answer #1

answer) we know the formula for acceleration of the electron is given by

a=qE/m=-1.6*10-19*(-1.6*10-10)jm/s2)/9.1*10-31=28.1m/s2 j

and the gravitational acceleration=-9.8m/s2 ( downwards)

so total acceleration=28.1-9.8=18.3m/s2

now we know the formula

v2=u2+2as=0+2*18.3*1.2*10-6m=43.92*10-6m/s

v=\sqrt{}43.92*10-6=6.63*10-3m/s=6.63mm/s

so the answer is 6.63mm/s or 6.6mm/s or 7mm/s